### v2

# Imbalance-Robust and Sampling-Efficient Continuous Conditional GANs via Adaptive Vicinity and Auxiliary Regularization

Abstract
Recent advances in conditional generative modeling have introduced Continuous conditional Generative Adversarial Network (CcGAN) and Continuous Conditional Diffusion Model (CCDM) for estimating high-dimensional data distributions conditioned on scalar, continuous regression labels (e.g., angles, ages, or temperatures). However, these approaches face fundamental limitations: CcGAN suffers from data imbalance due to fixed-size vicinity constraints, while CCDM requires computationally expensive iterative sampling. We present CcGAN-AVAR, an enhanced CcGAN framework that addresses both challenges: (1) leveraging the GAN framework’s native one-step generation to overcome CCDMs’ sampling bottleneck (achieving 300×-2000× faster inference), while (2) two novel components specifically target data imbalance - an adaptive vicinity mechanism that dynamically adjusts vicinity’s size, and a multi-task discriminator that constructs two regularization terms (through auxiliary regression and density ratio estimation) to significantly improve generator training. Extensive experiments on four benchmark datasets (64×64 to 192×192 resolution) across eight challenging imbalanced settings demonstrate that CcGAN-AVAR achieves state-of-the-art generation quality while maintaining sampling efficiency.
Codehttps://github.com/UBCDingXin/CcGAN-AVAR
1 Introduction
[…]
Figure 1: The illustration of CCGM with data imbalance in UTKFace and Steering Angle datasets.
To mitigate data scarcity, Ding et al. (2021, 2023b) proposed the Continuous conditional Generative Adversarial Network (CcGAN), which estimates the target data distribution $p_{r}(\bm{x}|y_{c})$ by incorporating training samples within a fixed-size hard or soft vicinity of the conditional label $y_{c}$. While these vicinities enable CcGAN to trade some label consistency—the degree to which generated samples align with the given condition—for improved training stability and generation quality, Figure 2(a)–(b) reveals a critical limitation: the fixed-size vicinities may unnecessarily compromise too much label consistency for densely sampled label values, ultimately leading to suboptimal generation performance.
As an alternative to CcGAN, Ding et al. (2025) recently integrated CcGAN techniques into diffusion models, proposing the Continuous Conditional Diffusion Model (CCDM). Although CCDM exhibits superior generation quality and robustness to data imbalance compared to CcGAN, its sampling process is 700–2000× slower, severely limiting practical deployment. While distillation-based one-step sampling mitigates this latency, it incurs significant generation quality degradation and additional training overhead (Ding et al. 2025). Thus, achieving both high-quality generation and efficient sampling remains an open challenge in CCGM.
Motivated by these limitations, we present CcGAN-AVAR, an improved CcGAN framework that achieves generation quality comparable to or surpassing CCDM while maintaining significantly faster sampling speeds (300×-2000× acceleration). Our key contributions include: (1) a novel soft/hybrid Adaptive Vicinity (AV) mechanism that overcomes the limitations of conventional hard/soft vicinities in handling data imbalance; (2) a multi-task discriminator with dedicated auxiliary branches for noise-insensitive regression and density ratio estimation, generating two auxiliary regularization terms that substantially improve generation quality and label consistency; (3) RC-49-I, an imbalanced variant of the RC-49 dataset (Ding et al. 2023b) incorporating three distinct imbalance patterns (unimodal, bimodal, and trimodal distributions) for thorough evaluation of CCGM methods under realistic data scarcity conditions; and (4) an optimized, unified codebase for CcGAN supporting multiple benchmark datasets, flexible network architecture and vicinity selection, various loss functions, exponential moving average, and mixed-precision training.
2 Related Work
2.1 Continuous Conditional Generative Models
To address data scarcity in CCGM tasks, CcGAN (Ding et al. 2021, 2023b) leverages samples within a label-dependent vicinity of $y_{c}$ to approximate the data distribution $p_{r}(\bm{x}|y_{c})$. The framework introduces two distinct vicinity formulations (hard vicinity and soft vicinity), which leads to the Hard/Soft Vicinal Discriminator Loss (HVDL/SVDL) and the corresponding generator loss for CcGAN:
$\mathcal{L}(D)=$
$-\frac{1}{N^{r}}\sum_{j=1}^{N^{r}}\sum_{i=1}^{N^{r}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\left[W^{r}_{i}\log(D(\bm{x}_{i}^{r},y_{j}^{r}+\epsilon))\right]$
$-\frac{1}{N^{g}}\sum_{j=1}^{N^{g}}\sum_{i=1}^{N^{g}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\left[W^{g}_{i}\log(1-D(\bm{x}_{i}^{g},y_{j}^{g}+\epsilon))\right],$ (1)
$\mathcal{L}(G)=$
$-\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\log(D(G(\bm{z}_{i},y_{i}^{g}+\epsilon),y_{i}^{g}+\epsilon)),$ (2)
where $D$ and $G$ denote the discriminator and generator networks, superscripts rr/$g$ indicate real/fake samples, $\bm{z}_{i}$ represents Gaussian noise, $\sigma$ is a positive hyperparameter, and $N^{r}$$N^{g}$ denote real/fake sample sizes. As illustrated in Figure 2(a), the hard vicinity defines a closed interval centered at $y_{c}$ with a fixed radius $\kappa$. Thus, the weights $W^{r}_{i}$ and $W^{g}_{i}$ for HVDL are defined as follows:
$W^{r}_{i}=\frac{\mathds{1}_{\{|y_{j}^{r}+\epsilon-y_{i}^{r}|\leq\kappa\}}}{N_{y^{r}_{j}+\epsilon,\kappa}^{r}},\quad W^{g}_{i}=\frac{\mathds{1}_{\{|y_{j}^{g}+\epsilon-y_{i}^{g}|\leq\kappa\}}}{N_{y^{g}_{j}+\epsilon,\kappa}^{g}},$ (3)
with $\mathds{1}_{\{\cdot\}}$ as the indicator function and $N_{y_{c},\kappa}$ counting samples in the hard vicinity $y_{c}-\kappa,y_{c}+\kappa]$. Ding et al. (2023b) provides a rule of thumb to compute $\kappa$ and $\sigma$ from samples. For soft vicinity, all samples are used to estimate $p_{r}(\bm{x}|y_{c})$, but each sample $\bm{x}_{i}$ is assigned with a weight $W_{i}$, with the following formulation:
$W^{r}_{i}=\frac{w(y_{i}^{r},y_{j}^{r}+\epsilon)}{\sum_{i=1}^{N^{r}}w(y_{i}^{r},y_{j}^{r}+\epsilon)},W^{g}_{i}=\frac{w(y_{i}^{g},y_{j}^{g}+\epsilon)}{\sum_{i=1}^{N^{g}}w(y_{i}^{g},y_{j}^{g}+\epsilon)}.$ (4)
As illustrated in Figure 2(b), the soft vicinity weights in Eq. (4) follow an exponential decay curve $w(y,y^{\prime})=e^{-\nu(y-y^{\prime})^{2}}$, where the decay rate $\nu>0$ controls the weighting sensitivity to label distances. Ding et al. (2023b) set $\nu=1/\kappa^{2}$ and exclude samples with soft weights below $10^{-3}$ from discriminator updates for computational efficiency.
Figure 2: Comparison of traditional fixed-size (hard/soft) vicinities versus our proposed (soft/hybrid) adaptive vicinities. Bar heights represent sample sizes per label value, with $y_{c}^{1}$ and $y_{c}^{2}$ marking two conditional label values exhibiting different local sample densities.
The conventional hard/soft vicinity (with constant $\kappa$ and $\nu$) often produces suboptimal generation quality on imbalanced datasets. As Figure 2(a)-(b) shows, while samples cluster densely around $y_{c}^{2}$, the fixed-size vicinity formulation either incorporates excessive samples (hard vicinity) or over-weights distant samples (soft vicinity), both of which significantly degrade label consistency.
Recent work by Ding et al. (2025) introduces CCDM, which incorporates vicinal techniques into diffusion models as a replacement for CcGAN. Although both their empirical study and the experiments in this paper (Tables 1 and 2) show CCDM outperforms conventional CcGANs while maintaining data imbalance robustness, its prohibitively slow sampling speed significantly limits practical applicability.
2.2 Sample-Based Density Ratio Estimation
Ding et al. (2023a) introduced cDR-RS, a sample-based Density Ratio Estimation (DRE) method for conditional GANs that computes $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$. The estimated density ratios enable rejection sampling to enhance output quality in both class-conditional GANs and CcGANs.
cDR-RS’s DRE framework employs a two-stage architecture: a fixed, pre-trained encoder that maps input images $\bm{x}$ into latent representations $\bm{h}$, and a trainable 5-layer fully-connected network $f_{\text{dre}}$ that processes $\bm{h}$ for density ratio estimation. During optimization, only $f_{\text{dre}}$ is updated using the following penalized Softplus loss:
$\mathcal{L}(f_{\text{dre}})=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left[\delta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-\eta(f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g}))\right]$
$-\frac{1}{N^{r}}\sum_{i=1}^{N_{r}}\delta(f_{\text{dre}}(\bm{h}_{i}^{r}|y_{i}^{r}))+\lambda_{\text{dre}}\cdot\left(\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}f_{\text{dre}}(\bm{h}_{i}^{g}|y_{i}^{g})-1\right)^{2},$ (5)
where $\delta$ and $\eta$ denote respectively the Sigmoid and Softplus functions, and $\lambda_{\text{dre}}$ is a hyperparameter for regularization.
3 Method
3.1 Overview
In this section, we present CcGAN-AVAR, an improved CcGAN framework designed to address data imbalance challenges. As illustrated in Figure 3, our approach introduces two key innovations: (1) a novel soft/hybrid adaptive vicinty mechanism that dynamically adjusts to local sample density, and (2) a multi-task discriminator that generates two auxiliary regularization terms to enhance generator training.
Figure 3: Overall workflow of CcGAN-AVAR
3.2 Adaptive Vicinity for Imbalanced Data
Let $Q^{r}=\{(\bm{x}^{r}_{i},y^{r}_{i})\}_{i=1}^{N^{r}}$ denote a training set of $N^{r}$ real image-label pairs, where the labels consist of $M$ distinct values $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ sorted in strictly increasing order ($y_{(i)}<y_{(i+1)}$ for i=1,,Mi=1,...,M). For $i$-th distinct label value $y_{(i)}$, we denote by $N_{i}$ its corresponding sample count. In severely imbalanced datasets, $N_{i}$ exhibits substantial variation across different $i$, as demonstrated in Figure 1(b)-(c).
To address data imbalance, we propose two distinct types of Adaptive Vicinity (AV): Soft AV and Hybrid AV. Both approaches replace the vicinity parameters $\kappa$ and $\nu$ in Eqs. (3) and (4) with a new hyperparameter $N_{\text{AV}}$, which determines the minimum number of effective samples per vicinity. Practical guidance for setting $N_{\text{AV}}$ is provided in the Appendix.
(1) Soft Adaptive Vicinity
Building on CcGAN’s empirically demonstrated superiority of soft vicinity over hard vicinity (Ding et al. 2021, 2023b), we first propose a soft AV mechanism (Figure 2(c)). During discriminator training, we create a target conditional label $y_{c}$ by randomly adding Gaussian noise to a training label (e.g., $y_{j}^{r}+\epsilon$ in Eq. (1)). Without loss of generality, assuming $y_{c}$ lies between $y_{(i-1)}$ and $y_{(i)}$, we initialize a hard vicinity centered at $y_{c}$ with zero-radius boundaries ($\kappa_{l}=0$, $\kappa_{r}=0$) containing $N_{c}=0$ samples. This vicinity then extends iteratively until $N_{c}\geq N_{\text{AV}}$: at each step, the vicinity extends leftwards if the leftmost external label is closer to $y_{c}$ than the rightmost external label, and rightwards otherwise. Following each extension, we update either $\kappa_{l}$ or $\kappa_{r}$ to reflect the distance of the newly incorporated label to $y_{c}$, while incrementing $N_{c}$ by the sample count at this label. Upon reaching $N_{c}\geq N_{\text{AV}}$, we obtain the final radius $\kappa_{y_{c}}=\max(\kappa_{l},\kappa_{r})$, which determines the local weight decay rate $\nu_{y_{c}}=1/\kappa_{y_{c}}^{2}$. Soft weights are then assigned to all training samples via Eq. (4) using $\nu_{y_{c}}$, with Figure 4 illustrating this process.
This adaptive approach automatically adjusts to local sample density: in dense regions (small $\kappa_{y_{c}}$), weights decay rapidly to maintain label consistency while preserving diversity; in sparse regions (large $\kappa_{y_{c}}$), weights decay slowly to incorporate more distant samples, ensuring training stability and preventing mode collapse.
Figure 4: Illustrative example: Building the soft or hybrid adaptive vicinity for condition $y_{c}$ through 3 extension steps
(2) Hybrid Adaptive Vicinity
We further propose a hybrid AV mechanism (Figs. 2(d) and 4) that extends the soft AV approach. This hybrid variant incorporates truncated weight decay by assigning zero weights to samples outside the interval $y_{c}-\kappa_{y_{c}},y_{c}+\kappa_{y_{c}}]$. By combining the adaptive weighting of soft AV with the boundary constraints of hard vicinity, this approach can further enhances label consistency.
3.3 Auxiliary Generator Regularization
(1) A Multi-Task Discriminator
We propose a multi-task discriminator framework $D(\bm{x},y)$, with architecture illustrated in Figure 3 (a). In our implementation, the backbone of $D$ consists of the convolutional blocks from SNGAN (Miyato et al. 2018), denoted as $D_{\text{conv}}$, though this component can be readily replaced with alternative GAN architectures such as DCGAN (Radford, Metz, and Chintala 2015), SAGAN (Zhang et al. 2019), BigGAN (Brock, Donahue, and Simonyan 2019), or GigaGAN (Kang et al. 2023). The convolutional backbone is followed by three parallel fully-connected branches: an one-layer adversarial branch $f_{\text{adv}}$, a two-layer regression branch $f_{\text{reg}}$, and a three-layer density ratio estimation branch $f_{\text{dre}}$.
The adversarial output, denoted by $D_{\text{adv}}(\bm{x},y)$, is generated by processing the extracted features $\bm{h}^{\bm{x}}=D_{\text{conv}}(\bm{x},y)$ through the adversarial branch $f_{\text{adv}}$ combined with label projection (Miyato et al. 2018). This output is then used to compute the adversarial loss $\mathcal{L}^{D}_{\text{adv}}$ according to Eqs. (1) and (4). Notably, the proposed soft or hybrid AV mechanism is fully compatible with this computation.
The regression branch $f_{\text{reg}}$ predicts the label of the input image through $\hat{y}=f_{\text{reg}}(D_{\text{conv}}(\bm{x},y))$. To optimize this branch, we employ a $\gamma$-insensitive hinge loss adapted from support vector regression (Murphy 2022), formulated as:
$\mathcal{L}^{D}_{\text{reg}}=$ $\frac{1}{N^{r}}\sum_{i=1}^{N^{r}}\mathds{E}_{\epsilon\sim\mathcal{N}(0,\sigma^{2})}\left[\max\left(|y_{i}^{r}+\epsilon-\hat{y}_{i}^{r}|-\gamma,0\right)\right]$
$+\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\max\left(|y_{i}^{g}-\hat{y}_{i}^{g}|-\gamma,0\right),$ (6)
where $y_{i}^{r}$ and $y_{i}^{g}$ denote the ground truth label for $\bm{x}_{i}^{r}$ and the conditional label used to generate $\bm{x}_{i}^{g}$, respectively, while $\hat{y}_{i}^{r}$ and $\hat{y}_{i}^{g}$ represent the corresponding prediction from $f_{\text{reg}}$. The elasticity parameter $\gamma\geq 0$, empirically set as the maximum raidus $\kappa_{y_{i}^{r}+\epsilon}$ in a training batch, accounts for Gaussian noise $\epsilon$ in real labels and helps mitigate misalignment between $\bm{x}_{i}^{g}$ and its conditional label $y_{i}^{g}$. In practice, we often replace $y_{i}^{g}$ with $\tilde{y}^{g}_{i}$, where $\tilde{y}^{g}_{i}$ denotes the output of a ResNet-18 (He et al. 2016) separately trained for label regression on $\bm{x}_{i}^{g}$.
The DRE branch $f_{\text{dre}}$ predicts the conditional density ratio $r(\bm{x}|y)=p_{r}(\bm{x}|y)/p_{g}(\bm{x}|y)$ for real and fake images. This branch is optimized through the loss function $\mathcal{L}_{\text{dre}}^{D}$ (Eq. (5)), where $\lambda_{\text{dre}}=10^{-2}$. Unlike Ding et al. (2023a), our feature extractor $D_{\text{conv}}$ is jointly optimized with $f_{\text{dre}}$.
The complete discriminator loss function is defined as:
$\mathcal{L}^{D}=\mathcal{L}^{D}_{\text{adv}}+\lambda^{D}_{\text{reg}}\cdot\mathcal{L}^{D}_{\text{reg}}+\lambda^{D}_{\text{dre}}\cdot\mathcal{L}^{D}_{\text{dre}},$ (7)
where $\lambda^{D}_{\text{reg}}\geq 0$ and $\lambda^{D}_{\text{dre}}\geq 0$ are the weighting coefficients.
(2) Auxiliary Regularization for Generator Training
The generator’s adversarial loss $\mathcal{L}^{G}_{\text{adv}}$ (Eq. (2)) is supplemented by two additional regularization terms derived from the multi-task discriminator: a regression penalty and a $f$-divergence penalty. The regression penalty utilizes mean absolute error to quantify label consistency:
$\mathcal{L}^{G}_{\text{reg}}=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}|y_{i}^{g}-\hat{y}_{i}^{g}|.$ (8)
This formulation directly minimizes the discrepancy between $\bm{x}_{i}^{g}$ and $\hat{y}_{i}^{g}$, thereby enforcing stronger label alignment during image synthesis. The $f$-divergence penalty, implemented through Pearson $\chi^{2}$ divergence with $f=(t-1)^{2}$ (Nowozin, Cseke, and Tomioka 2016), provides distributional matching:
$\mathcal{L}^{G}_{f}=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left(f_{\text{dre}}(D_{\text{conv}}(\bm{x}_{i}^{g},y_{i}^{g}))-1\right)^{2}.$ (9)
This term estimates the $f$-divergence between $p_{r}(\bm{x}|y)$ and $p_{g}(\bm{x}|y)$, guiding the generator towards more realistic data synthesis. The complete generator loss function combines these components through weighted summation:
$\mathcal{L}^{G}=\mathcal{L}^{G}_{\text{adv}}+\lambda^{G}_{\text{reg}}\cdot\mathcal{L}^{G}_{\text{reg}}+\lambda^{G}_{f}\cdot\mathcal{L}^{G}_{f},$ (10)
where $\lambda^{G}_{\text{reg}}\geq 0$ and $\lambda^{G}_{f}\geq 0$ are the weighting coefficients.
4 Experiments
4.1 Experimental Setup
Datasets. Building upon established experimental protocols (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we conduct comprehensive evaluations across multiple benchmark datasets: RC-49 (Ding et al. 2023b) (44,051 chair images at $64\times 64$ resolution with rotation angles from $0.1^{\circ}$ to $89.9\circ$ in $0.1^{\circ}$ increments), UTKFace (Zhang, Song, and Qi 2017) (14,723 facial images across $64\times 64$, $128\times 128$, and $192\times 192$ resolutions with age labels from 1 to 60), and Steering Angle (Chen 2018) (12,271 driving images at $64\times 64$ and $128\times 128$ resolutions covering steering angles from $80^{\circ}$ to $80^{\circ}$). For UTKFace and Steering Angle experiments, all available samples are used for both training and evaluation. The RC-49 experiments employ a curated balanced subset of 11,250 images (25 samples each for 450 distinct angle labels) for training while utilizing the complete collection of 44,051 images for comprehensive evaluation.
To rigorously evaluate model robustness to data imbalance, we introduce RC-49-I—a specialized variant of RC-49 comprising three subsets with distinct label distribution patterns (unimodal, bimodal, and trimodal; see Figure 5). This framework enables systematic comparison between three challenging imbalanced scenarios (UTKFace, Steering Angle, and RC-49-I) and the balanced RC-49 baseline. All models trained on RC-49-I are evaluated using the complete RC-49 dataset (44,051 samples) to ensure consistent assessment.
Figure 5: Three data imbalance patterns in RC-49-I
Compared Methods. Our comparative evaluation examines six continuous conditional generative models: CcDPM (Zhao et al. 2024) and CCDM (Ding et al. 2025) as diffusion-based approaches; CcGAN (Ding et al. 2023b) and Dual-NDA (Ding, Wang, and Xu 2024) representing vanilla CcGAN architectures; along with our proposed variants CcGAN-AVAR-S (soft adaptive vicinity) and CcGAN-AVAR-H (hybrid adaptive vicinity).
Implementation Setup. We utilize the released results of CcDPM, CCDM, CcGAN, and Dual-NDA on RC-49, UTKFace, and Steering Angle from (Ding et al. 2025). For our newly introduced RC-49-I dataset, we implement these baseline methods while maintaining identical experimental configurations to those used for RC-49. As a key contribution of this work, we develop and release an optimized, unified codebase for CcGAN and our proposed CcGAN-AVAR framework, featuring support for multiple benchmark datasets, flexible backbone network selection, diverse vicinity selection strategies, various GAN losses and regularization penalties, exponential moving average (Karras et al. 2024), and mixed-precision training (Micikevicius et al. 2018).
For CcGAN-AVAR-S and CcGAN-AVAR-H implementations, we set $N_{\text{AV}}$ to 50 (RC-49), 30 (RC-49-I), 400 (UTKFace), and 20 (Steering Angle). All experiments utilize a SNGAN backbone architecture with fixed hyperparameters $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$ and $\lambda_{f}^{G}=0.5$. The parameter $\lambda_{\text{dre}}^{D}$ is set to 1.0 for both variants on RC-49 and for CcGAN-AVAR-H on RC-49-I, while using 0.5 in all other cases. To compute the $\gamma$-insensitive hinge loss for the regression branch of $D$, we pre-train a ResNet-18 model on each dataset to predict labels from images. Complete training configurations are provided in the Appendix and our codebase.
Evaluation Setup. For systematic evaluation, each trained model generates a predetermined number of images at specified evaluation centers across all datasets. On RC-49 and RC-49-I, models produce 200 fake images for each of 899 rotation angles spanning $0.1^{\circ}$ to $89.9$ in $0.1^{\circ}$ increments. For UTKFace, models generate 1,000 images per age from 1 to 60 years. The Steering Angle evaluation employs 2,000 uniformly distributed points across $80^{\circ},80^{\circ}]$, with 50 images synthesized per point. This protocol yields total generation counts of 179,800 images (RC-49 and RC-49-I), 60,000 images (UTKFace), and 100,000 images (Steering Angle) for each evaluated method.
Following the evaluation protocol established in (Ding et al. 2023b; Ding, Wang, and Xu 2024; Ding et al. 2025), we assess generated images using one primary metric and three complementary metrics. The Sliding Fréchet Inception Distance (SFID) (Ding et al. 2023b) serves as our primary evaluation metric, computed as the averaged FID score across all evaluation centers. Additionally, we employ three complementary metrics: (1) the Naturalness Image Quality Evaluator (NIQE) (Mittal, Soundararajan, and Bovik 2012) for assessing visual realism, (2) Diversity (Ding et al. 2023b) for quantifying sample variety, and (3) Label Score (Ding et al. 2023b) for measuring label consistency. For SFID, NIQE, and Label Score, lower values indicate superior performance, whereas higher Diversity scores reflect greater sample variety. We refer interested readers to Ding et al. (2025) for complete technical details regarding these evaluation metrics.
Furthermore, we evaluate the sampling efficiency of each compared model by reporting its sampling speed (images/second) and GPU memory consumption (GiB), measured on a Linux system equipped with an NVIDIA RTX 4090D.
4.2 Experimental Results
We conduct a comprehensive evaluation of candidate methods across nine distinct settings, including eight imbalanced configurations and one balanced setting. The empirical results, presented in Tables 1 and 2 along with Figure 6, reveal several key findings:
  • The comparative analysis between RC-49 and RC-49-I reveals significant performance degradation for both vanilla CcGAN and its Dual-NDA variant when applied to the imbalanced RC-49-I dataset, demonstrating the substantial impact of data imbalance on CcGANs.
  • CCDM is more robust than CcGANs with a small drop of generation performance on imbalanced datasets. However, the iterative sampling process makes CCDM 700-2000× slower at sampling than CcGAN/Dual-NDA.
  • The two CcGAN-AVAR variants achieve superior performance across most settings by SFID metrics, maintaining comparable or better sample diversity than CcGAN/Dual-NDA while significantly improving label consistency - confirming the efficacy of their adaptive vicinity and auxiliary regularization designs. Crucially, CcGAN-AVAR preserves consistent generation quality between balanced (RC-49) and imbalanced (RC-49-I) datasets, unlike baseline methods that exhibit significant degradation, demonstrating exceptional robustness to data skews. Furthermore, the hybrid AV variant typically outperforms soft AV in label consistency across most configurations.
  • CcGAN-AVAR achieves 300×-2000× faster generation than CCDM while maintaining sampling speeds comparable to CcGAN and Dual-NDA, along with moderate GPU memory requirements.
Dataset
Method
SFID
$\downarrow$
NIQE $\downarrow$ Diversity $\uparrow$
Label
Score
$\downarrow$
Speed
(img/sec)
$\uparrow$
Memory
$\downarrow$
RC-49
($\times64$)
CcDPM (AAAI’24) 0.970 2.153 3.581 24.174 2.92 5.07
CCDM (arXiv’25) 0.049 2.086 3.698 1.074 2.92 5.53
CcGAN (T-PAMI’23) 0.126 1.809 3.451 2.655 6243.06 2.62
Dual-NDA (AAAI’24) 0.148 1.808 3.344 2.211 6243.06 2.62
CcGAN-AVAR-S 0.048 1.728 3.705 1.273 4162.04 3.71
CcGAN-AVAR-H 0.042 1.734 3.723 1.270 4162.04 3.71
RC-49-I
($\times64$)
CcDPM (AAAI’24) 0.925 2.136 3.585 23.749 2.92 5.07
CCDM (arXiv’25) 0.064 2.023 3.642 1.417 2.92 5.53
CcGAN (T-PAMI’23) 0.151 1.880 3.400 2.487 6243.06 2.62
Dual-NDA (AAAI’24) 0.192 1.941 3.223 2.436 6243.06 2.62
CcGAN-AVAR-S 0.054 1.729 3.704 1.504 4162.04 3.71
CcGAN-AVAR-H 0.050 1.794 3.711 1.425 4162.04 3.71
RC-49-I
($\times64$)
CcDPM (AAAI’24) 0.850 2.214 3.572 21.972 2.92 5.07
CCDM (arXiv’25) 0.066 2.012 3.659 1.272 2.92 5.53
CcGAN (T-PAMI’23) 0.205 1.906 3.293 2.825 6243.06 2.62
Dual-NDA (AAAI’24) 0.235 1.886 3.162 2.317 6243.06 2.62
CcGAN-AVAR-S 0.052 1.736 3.716 1.446 4162.04 3.71
CcGAN-AVAR-H 0.050 1.729 3.706 1.402 4162.04 3.71
RC-49-I
($\times64$)
CcDPM (AAAI’24) 0.940 2.224 3.551 23.902 2.92 5.07
CCDM (arXiv’25) 0.060 2.047 3.664 1.117 2.92 5.07
CcGAN (T-PAMI’23) 0.171 1.873 3.296 2.813 6243.06 2.62
Dual-NDA (AAAI’24) 0.168 1.949 3.302 2.575 6243.06 2.62
CcGAN-AVAR-S 0.049 1.735 3.716 1.490 4162.04 3.71
CcGAN-AVAR-H 0.051 1.729 3.707 1.462 4162.04 3.71
UTKFace
($\times64$)
CcDPM (AAAI’24) 0.466 1.560 1.211 6.868 2.10 5.70
CCDM (arXiv’25) 0.363 1.542 1.184 6.164 2.10 6.14
CcGAN (T-PAMI’23) 0.413 1.733 1.329 8.240 5555.56 2.44
Dual-NDA (AAAI’24) 0.396 1.678 1.298 6.765 5555.56 2.44
CcGAN-AVAR-S 0.358 1.703 1.298 7.222 4166.67 3.71
CcGAN-AVAR-H 0.356 1.691 1.278 6.696 4166.67 3.71
Steering Angle
($\times64$)
CcDPM (AAAI’24) 0.939 1.761 1.150 10.999 3.33 2.28
CCDM (arXiv’25) 0.742 1.778 1.088 5.823 3.33 2.74
CcGAN (T-PAMI’23) 1.334 1.784 1.234 14.807 3086.42 2.50
Dual-NDA (AAAI’24) 1.114 1.738 1.251 11.809 3086.42 2.50
CcGAN-AVAR-S 0.748 1.810 1.176 6.408 1157.41 3.86
CcGAN-AVAR-H 0.809 1.800 1.204 6.963 1157.41 3.86
Table 1: Comparison of generation quality and sampling efficiency at $64\times 64$ resolution. Performance is assessed using four metrics: SFID (overall quality), NIQE (visual realism), Diversity (sample diversity), and Label Score (label consistency), along with sampling speed and GPU memory consumption measured on an RTX 4090D. $\downarrow$” ($\uparrow$”) indicates lower (higher) values are preferred. The best and second-best results are highlighted in bold and underlined, respectively.
Dataset
Method
SFID
$\downarrow$
NIQE $\downarrow$ Diversity $\uparrow$
Label
Score
$\downarrow$
Speed
(img/sec)
$\uparrow$
Memory
$\downarrow$
UTKFace
($\times128$)
CcDPM (AAAI’24) 0.529 1.114 1.195 7.933 1.12 17.91
CCDM (arXiv’25) 0.319 1.077 1.178 6.359 1.12 19.52
CcGAN (T-PAMI’23) 0.367 1.113 1.199 7.747 1282.05 7.58
Dual-NDA (AAAI’24) 0.361 1.081 1.257 6.310 1282.05 7.58
CcGAN-AVAR-S 0.299 1.191 1.223 6.622 1190.48 11.57
CcGAN-AVAR-H 0.297 1.173 1.251 6.586 1190.48 11.57
Steering Angle
($\times128$)
CcDPM (AAAI’24) 1.285 1.989 1.203 18.325 1.23 5.27
CCDM (arXiv’25) 0.987 1.977 1.118 11.829 1.23 6.85
CcGAN (T-PAMI’23) 1.689 2.411 1.088 18.438 1543.21 7.58
Dual-NDA (AAAI’24) 1.390 2.135 1.133 14.099 1543.21 7.58
CcGAN-AVAR-S 0.803 2.350 1.079 6.908 1207.73 11.57
CcGAN-AVAR-H 0.888 2.288 1.123 7.507 1207.73 11.57
UTKFace
($\times192$)
CcDPM (AAAI’24) 0.970 1.522 1.187 11.224 0.80 19.93
CCDM (arXiv’25) 0.467 1.242 1.148 7.336 0.80 22.99
CcGAN (T-PAMI’23) 0.499 1.661 1.207 7.885 617.28 18.18
Dual-NDA (AAAI’24) 0.487 1.483 1.201 6.730 617.28 18.18
CcGAN-AVAR-S 0.455 1.621 1.104 6.433 1111.11 13.18
CcGAN-AVAR-H 0.435 1.588 1.187 6.377 1111.11 13.18
Table 2: Comparison of generation quality and sampling efficiency at high resolutions ($64\times 64$) across methods. The best and second-best results are highlighted in bold and underlined, respectively.
Figure 6: Visual comparison of FID, NIQE, Diversity, and Label Score versus angle (evaluation center) for three compared methods on RC-49-I (bimodal).
4.3 Ablation Study
We perform ablation studies on RC-49-I and UTKFace at 64×64 resolution to systematically evaluate the impact of individual components and CcGAN-AVAR’s key hyperparameter, as outlined blow.
  • The Effects of AV and Regularization: Our component-wise analysis of CcGAN-AVAR on RC-49-I and UTKFace (Table 3) reveals that both soft and hybrid adaptive vicinity mechanisms substantially improve label consistency while maintaining or enhancing sample diversity. The auxiliary regularization terms further boost label consistency, resulting in significant overall quality improvements. It is worth noting that even the baseline configuration (without adaptive vicinity or auxiliary regularization) outperforms CcGAN/Dual-NDA (Table 1), due to a deeper GAN network and more advanced training techniques such as exponential moving average.
  • The Effects of $N_{\text{AV}}$: We investigate the impact of the key hyperparameter $N_{\text{AV}}$ on CcGAN-AVAR-H’s performance (Figure 7), conducting this ablation study with auxiliary regularization disabled to isolate AV’s effect. The results demonstrate substantial performance degradation when either disabling AV entirely ($N_{\text{AV}}=0$) or employing an excessively wide vicinity (large $N_{\text{AV}}$ values), while moderate $N_{\text{AV}}$ values consistently yield acceptable results. The dashed green lines in Figure 7 correspond to the optimal values selected for our main experiments.
Dataset Configuration
SFID
NIQE Diversity
Label
Score
Soft AV Hybrid AV Aux. Reg. Aux. DRE
RC-49-I
($64\times 64$)
0.062 1.740 3.704 2.291
0.059 1.732 3.700 2.238
0.055 1.727 3.695 1.737
0.052 1.736 3.716 1.446
RC-49-I
($64\times 64$)
0.062 1.740 3.704 2.291
0.058 1.739 3.708 2.052
0.050 1.723 3.715 1.644
0.050 1.729 3.706 1.402
UTKFace
($64\times 64$)
0.447 1.738 1.243 8.274
0.442 1.750 1.305 7.867
0.432 1.747 1.307 8.073
0.358 1.703 1.298 7.222
UTKFace
($64\times 64$)
0.447 1.738 1.243 8.274
0.371 1.696 1.310 7.246
0.357 1.709 1.269 6.529
0.356 1.691 1.278 6.696
Table 3: Ablation study: Component-wise analysis of CcGAN-AVAR
Figure 7: Ablation study: SFID/Label Score versus $N_{\text{AV}}$
5 Conclusion
This paper presents CcGAN-AVAR, an advanced CcGAN framework that achieves state-of-the-art generation performance while maintaining exceptional sampling efficiency. The proposed framework introduces two key innovations: (1) soft and hybrid adaptive vicinity mechanisms that effectively address the severe data imbalance inherent in regression datasets, replacing conventional fixed-size vicinity approaches; and (2) a dual regularization strategy incorporating both auxiliary regression and $f$-divergence penalties during generator training, significantly enhancing output quality. Extensive experiments across eight challenging imbalanced settings with varying resolutions demonstrate CcGAN-AVAR’s consistent superiority over existing methods.
6 Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62306147, 62201404, U24A20326), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241901), Jiangsu Supplementary Funds of National Major Talent Project (2022-2024) (Grant No. R2024PT03), Nanjing Selective Funding Initiative for Returnee Innovation Projects (Grant No. R2024LZ02), the Startup Foundation for Introducing Talent of NUIST (Grant No. 2024r056), Yangtze River Delta Community of Sci-Tech Innovation (Grant No. 2023CSJZN0301), and Pujiang Talent Program (Grant No. 23PJ1412100).
Supplemental Material
Appendix A Adaptive Vicinity Implementation
A.1 Algorithm for Adaptive Vicinity Construction
Constructing the proposed soft or hybrid vicinity hinges on determining two critical parameters: the vicinity radius $\kappa_{y_{c}}$ and the weight decay rate $v_{y_{c}}$. A detailed procedure for computing these parameters is presented in Algorithm 1.
Data: Distinct training labels $Y^{u}=\{y_{(1)},...,y_{(M)}\}$; Sample count for each distinct training label $\{N_{1},...,N_{M}\}$;
Input: Conditional label $y_{c}$; the minimum effective sample count $N_{\text{AV}}$;
Result: Vicinity width $\kappa_{y_{c}}$ and weight decay rate $v_{y_{c}}$
1 $\kappa_{l}\leftarrow 0$, $\kappa_{r}\leftarrow 0$;
2 $N_{c}=0$;
3 $t_{l}=i-1$, $t_{r}=i$;
// Vicinity extension
4 while $N_{c}<N_{\text{AV}}$ do
5 if $|y_{(t_{l})}-y_{c}|<|y_{(t_{r})}-y_{c}|$ then
    // Extend leftwards
6    $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
7    $N_{c}\leftarrow N_{c}+N_{t_{l}}$;
8    $t_{l}\leftarrow t_{l}-1$;
9
10 else if $|y_{(t_{l})}-y_{c}|>|y_{(t_{r})}-y_{c}|$ then
    // Extend rightwards
11    $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
12    $N_{c}\leftarrow N_{c}+N_{t_{r}}$;
13    $t_{r}\leftarrow t_{r}+1$;
14
15 else
    // Extend in both directions
16    $\kappa_{l}\leftarrow|y_{(t_{l})}-y_{c}|$;
17    $\kappa_{r}\leftarrow|y_{(t_{r})}-y_{c}|$;
18    $N_{c}\leftarrow N_{c}+N_{t_{l}}+N_{t_{r}}$;
19    $t_{l}\leftarrow t_{l}-1$;
20    $t_{r}\leftarrow t_{r}+1$;
21
22   end if
23
24 end while
[…]
26 $\nu_{y_{c}}\leftarrow 1/\kappa_{y_{c}}$;
Algorithm 1 Soft/Hybrid Adaptive Vicinity Construction for Conditional Label $y_{c}$. Without loss of generality, we assume $y_{c}$ lies between $y_{(i-1)}$ and $y_{(i)}$.
A.2 Heuristic Selection of $N_{c}$
The minimum number of effective samples per vicinity ($N_{c}$) serves as the crucial hyperparameter in CcGAN-AVAR, with ablation studies on RC-49-I and UTKFace demonstrating optimal performance when $N_{c}$ is properly calibrated within a moderate range, as evidenced in Figure S.8.
To determine $N_{c}$ in practice, we propose a heuristic approach:
  • First, identify all the distinct training labels
    $Y^{u}=\{y_{(1)},...,y_{(M)}\}$ (S.11)
    and their sample counts
    $\mathcal{N}=\{N_{1},...,N_{M}\}.$ (S.12)
  • Then, compute the summed counts for each and its immediate right neighbor, with results stored in
    $\mathcal{S}=\{S_{1},...,S_{M-1}\}.$ (S.13)
  • Finally, compute the averaged sample count $\bar{S}=\frac{1}{M-1}\sum_{i=1}^{M-1}S_{i}$, and set $N_{c}$ to an integer near $\bar{S}$ with potential minor adjustments based on the sparsity characteristics of the data distribution.
For instance, on the bimodal RC-49-I dataset where $\bar{S}\approx 25.279$, we set $N_{c}=30$. For UTKFace with $\bar{S}\approx 480$, we chose $N_{c}=400$. In the case of Steering Angle, where many $S_{i}\in\mathcal{S}$ cluster around 2 despite $\bar{S}\approx 13$, we selected $N_{c}=20$ to mitigate the impact of sparse label regions.
Figure S.8: Ablation analysis of adaptive vicinity size ($N_{c}$) effects on CcGAN-AVAR-H performance, comparing SFID and Label Score metrics across RC-49-I and UTKFace datasets at 64×64 resolution. Dashed green lines correspond to the parameter values used in the primary experiment. Auxiliary regularization was disabled to isolate the contribution of adaptive vicinity.
Appendix B Auxiliary Regularization Implementation
B.1 Network Architecture
CcGAN-AVAR adopts a SNGAN-based architecture (Miyato et al. 2018), incorporating an additional residual block in the generator compared to the original implementation in (Ding et al. 2023b). While Ding et al. (2023b) recommends SAGAN (Zhang et al. 2019) for high-resolution scenarios, we observe that its self-attention mechanism exhibits limited compatibility with exponential moving average (EMA) (Karras et al. 2024) – a technique that effectively improves SNGAN performance in our experiments.
A key contribution of this work is our modified discriminator architecture, which integrates both a two-layer regression branch and a three-layer density ratio estimation (DRE) branch, as illustrated in Figures S.9 and S.10. The regression branch employs two linear layers with spectral normalization (SN) (Miyato et al. 2018) and ReLU activation, while the DRE branch replaces SN with group normalization (GN) (Wu and He 2020) using 8 groups. For high-resolution UTKFace (192×192) experiments, we incorporate a dropout layer after ReLU activation - crucially, not for the conventional purpose of preventing overfitting, but to address a distinct training challenge where the DRE branch otherwise fails to converge (evidenced by stagnant training loss).
Complete architectural details and implementations are provided in our code repository.
Figure S.9: Architecture of the two-layer fully connected regression branch $f_{\text{reg}}$ in the discriminator.
Figure S.10: Architecture of the three-layer fully connected DRE branch $f_{\text{dre}}$ in the discriminator, where Dropout is optional.
B.2 Relation to ACGAN
CcGAN-AVAR differs fundamentally from ACGAN (Odena, Olah, and Shlens 2017) in three key aspects. First, while ACGAN uses an auxiliary classification branch for class-conditional generation, its discriminator does not directly receive class information. In contrast, CcGAN-AVAR integrates a regression branch for continuous label prediction and explicitly injects condition labels into the discriminator via label projection (Figure 3). Second, whereas ACGAN relies on a cross-entropy loss for classification, CcGAN-AVAR employs a regression loss—specifically, a custom $\gamma$-insensitive hinge loss (Eq. (6)) designed to improve robustness against label noise, rather than a standard mean squared error. Third, and most distinctively, CcGAN-AVAR implements a density ratio estimation branch that imposes an $f$-divergence penalty during generator training - a feature entirely absent in the standard ACGAN architecture.
B.3 Extra Illustration for $f$-Divergence Penalty
The $f$-divergence (Nowozin, Cseke, and Tomioka 2016) between $p_{r}(\bm{x}|y)$ and $p_{g}(\bm{x}|y)$ is defined as:
$\text{Div}_{f}(p_{r}\|p_{g})$ $=\int_{\mathcal{X}}p_{g}(\bm{x}|y)f\left(\frac{p_{r}(\bm{x}|y)}{p_{g}(\bm{x}|y)}\right)\mathrm{d}\bm{x}$
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[f\left(\frac{p_{r}(\bm{x}|y)}{p_{g}(\bm{x}|y)}\right)\right],$ (S.14)
where $f$ is a convex function with $f(1)=0$. If $f$ is a quadratic function of the form $f(t)=(t-1)^{2}$, the $f$-divergence reduces to the Pearson $\chi^{2}$ divergence:
$\text{Div}_{\chi^{2}}(p_{r}\|p_{g})$ $=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[\left(\frac{p_{r}(\bm{x}|y)}{p_{g}(\bm{x}|y)}-1\right)^{2}\right]$
$=\mathbb{E}_{\bm{x}\sim p_{g}(\bm{x}|y)}\left[(r(\bm{x}|y)-1)^{2}\right].$ (S.15)
The proposed $f$-divergence penalty of CcGAN-AVAR is defined as follows:
$\mathcal{L}^{G}_{f}=\frac{1}{N^{g}}\sum_{i=1}^{N^{g}}\left(f_{\text{dre}}(D_{\text{conv}}(\bm{x}_{i}^{g},y_{i}^{g}))-1\right)^{2}.$ (S.16)
In Eq. (S.16), $f_{\text{dre}}(D_{\text{conv}}(\bm{x}_{i}^{g},y_{i}^{g}))$ estimates the conditional density ratio $r(\bm{x}|y)$. Thus, Eq. (S.16) approximates Eq. (S.15), and minimizing $\mathcal{L}^{G}_{f}$ encourages the generator to better match the target data distribution.
Appendix C Experimental Setup and Implementation
C.1 How RC-49-I is made?
Since RC-49 is a relatively simple synthetic dataset with a sufficiently large sample size for reliable evaluation, we construct three imbalanced subsets to thoroughly assess the robustness of CcGAN-AVAR and other baseline methods. As shown in Figure S.11, these subsets exhibit unimodal, bimodal, and trimodal label distributions. Using the unimodal case as an example, we describe the derivation procedure from RC-49 as follows:
  • First, we randomly select a distinct label $y_{m}\in Y^{u}$ as the mode of the imbalanced distribution.
  • Second, for each element $y_{(i)}$ in $Y^{u}$ excluding $y_{m}$, compute its absolute distance to $y_{m}$ as $d_{i}=|y_{(i)}-y_{m}|$.
  • Third, determine the mean sample count for $y_{(i)}$ using $\bar{N}^{\circ}_{i}=\max(1,\operatorname{int}(49\times e^{-\pi\cdot d_{i}}))$, where $\pi$ is a decay rate (set to 0.1 for unimodal distributions). The final sample count for $y_{(i)}$ is then calculate with noise perturbation: $N^{\circ}_{i}=\min(49,\max(0,\operatorname{int}(\bar{N}^{\circ}_{i}+\mathcal{N}(0,5))))$
  • Finally, we construct RC-49-I (unimodal) in a label-by-label manner by randomly selecting $N^{\circ}_{i}$ images from the complete set of 49 images for each $y_{(i)}$.
The bimodal and trimodal subsets require a more sophisticated generation procedure, with full implementation details available in our codebase.
Figure S.11: Three data imbalance patterns in RC-49-I
C.2 Hardware and Software Details
All experiments were conducted on NVIDIA RTX 4090D servers with the following hardware configurations:
  • $64\times 64$ resolution: 24GB GPU memory
  • $128\times 128$ and $192\times 192$ resolutions: 48GB GPU memory (RTX 4090D-48G)
The key software requirements include:
  • Operating System: Ubuntu 22.04 LTS
  • GPU Computing: CUDA 12.8
  • Core Software: Python 3.12.7; MATLAB R2021b
  • Python Packages: torch 2.7.0; torchvision 0.22.0; accelerate 1.6.0; numpy 1.26.4; scipy 1.13.1; h5py 3.11.0; Pillow 10.4.0; matplotlib 3.9.2
C.3 Training Configuration Details
For CcGAN-AVAR implementation, we train a ResNet-18 model from scratch on each dataset’s training set to predict image regression labels, using 200 epochs with a batch size of 256, base learning rate of 0.01, weight decay of $10^{-4}$, and enabled mixed-precision training.
Table S.4 details our CcGAN-AVAR training configurations, which employ a modified SNGAN backbone with exponential moving average and mixed-precision training - improvements over Ding et al. (2023b) [^1] https://github.com/UBCDingXin/improved˙CcGAN and Ding, Wang, and Xu (2024) [^2] https://github.com/UBCDingXin/Dual-NDA. Through mixed-precision training, our implementation demonstrates a 2× improvement in training speed and 50% reduction in memory usage compared to Ding et al. (2023b) when using identical architecture and batch size configurations.
Dataset Configuration
RC-49
[…]
num_D_steps=2, $N_{\text{AV}}=50$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=1$, $\lambda_{f}^{G}=0.5$
RC-49-I
[…]
num_D_steps=2, $N_{\text{AV}}=50$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=1$ (hybrid), $\lambda_{\text{dre}}^{G}=0.5$ (Soft), $\lambda_{f}^{G}=0.5$
RC-49-I
[…]
num_D_steps=2, $N_{\text{AV}}=50$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=1$ (hybrid), $\lambda_{\text{dre}}^{G}=0.5$ (Soft), $\lambda_{f}^{G}=0.5$
RC-49-I
[…]
num_D_steps=2, $N_{\text{AV}}=50$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=1$ (hybrid), $\lambda_{\text{dre}}^{G}=0.5$ (Soft), $\lambda_{f}^{G}=0.5$
UTKFace
($\times64$)
[…]
num_D_steps=2, $N_{\text{AV}}=400$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=0.5$, $\lambda_{f}^{G}=0.5$
Steering Angle
($\times64$)
SNGAN, gene_ch=64, disc_ch=64, hinge loss, DiffAugment, use EMA, steps=200K (hybrid) / 190K (soft),
[…]
UTKFace
($\times128$)
[…]
num_D_steps=2, $N_{\text{AV}}=400$, $\lambda_{\text{dre}}=10^{-2}$, $\lambda_{\text{reg}}^{D}=\lambda_{\text{reg}}^{G}=1$, $\lambda_{\text{dre}}^{G}=0.5$, $\lambda_{f}^{G}=0.5$
Steering Angle
($\times128$)
SNGAN, gene_ch=64, disc_ch=48, hinge loss, DiffAugment, use EMA, steps=200K (hybrid) / 190K (soft),
[…]
UTKFace
($\times192$)
SNGAN, gene_ch=64, disc_ch=48, hinge loss, DiffAugment, use EMA, steps=70K (hybrid) / 50K (soft),
[…]
DRE branch uses Dropout
Table S.4: Training Configurations for CcGAN-AVAR
C.4 Evaluation Configuration Details
Our evaluation protocol follows Ding et al. (2023b); Ding, Wang, and Xu (2024); Ding et al. (2025), with detailed specifications provided in Appendix S.V.B of Ding et al. (2025). For the RC-49-I experiments, we maintain identical evaluation procedures to those used for RC-49.
Appendix D Extra Experimental Results
Example synthetic images generated by Dual-NDA and CcGAN-AVAR for the UTKFace and Steering Angle experiments ($\times128$ resolution) are presented in Figs. S.12 and S.13.
Figure S.12: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the UTKFace ($\times128$) experiment.
Figure S.13: Some example fake images from Dual-NDA and CcGAN-AVAR-H in the Steering Angle ($\times128$) experiment.